OpenStudy (anonymous):

Solve for x: log x + log (x + 48) = 2

OpenStudy (anonymous):

log(x^2 + 48x) = 2 Just log means its base is 10, which is why logbase10(x^2+48x) = 2 Converting it into exponential form x^2 + 48x = 10^2 x^2 + 48x = 100 x^2 + 48x - 100 = 0 (x + 50)(x - 2) = 0 x = -50, 2

OpenStudy (anonymous):

Okay :) thanks! it kinda makes sense. Im not good at math, but you did put it out easily. Im alittle confused tho on y theres two answers.

OpenStudy (anonymous):

Oh! It can only be 2, because you can't have a negative number in logarithms, ex. log(-50). Sorry I forgot that.

OpenStudy (anonymous):

ah thanks! :) yes i see now.